OpenStudy (anonymous):

How much would $100 invested at 8% interest compounded continuously be worth after 15 years? Round your answer to the nearest cent. A. $220.00 B. $317.22 C. $332.01 D. $285.67

OpenStudy (kropot72):

\[\large A=Pe^{rt}=100e^{(0.08\times15)}\] Can you finish it now?

OpenStudy (anonymous):

yes i got C

OpenStudy (kropot72):

Good work! You are correct.

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (mathmath333):

@kropot72 but this formula gave this value \(\large\tt \begin{align} \color{black}{FV=PV(1+r)^n\\ =100(1+0.08)^{15}\\ =317.21}\end{align}\)

OpenStudy (kropot72):

@mathmath333 The reason for your result is that you have not used the formula for continuous compounding. There is a good explanation here: http://en.wikipedia.org/wiki/Compound_interest

OpenStudy (mathmath333):

oh,thnks

OpenStudy (kropot72):

You're welcome :)